OpenStudy (anonymous):

Help Please? A polynomial function can be written as (x - 2)(x - 3)(x + 5). What are the x-intercepts of the graph of this function? (2,0), (3,0), (-5,0) (-2,0),(-3,0),(5,0) (2,0),(3,0),(5,0) (-2,0),(-3,0),(-5,0)

OpenStudy (a1234):

Do (x - 2)(x - 3)(x + 5) = 0

OpenStudy (anonymous):

What? I have no idea what any of this means.

OpenStudy (anonymous):

2 , 3 and -5 so A

OpenStudy (a1234):

You're supposed to set each of those to 0. x - 2 = 0 x - 3 = 0 x + 5 = 0 Have you read your text?

OpenStudy (anonymous):

loli want an explanation?

OpenStudy (here_to_help15):

Hello @lolibear still need hhelp?

OpenStudy (anonymous):

well I will lol cause why not... So x - intercept means what are points where x value is zero. So if one of those parentheses equals 0 then we know it is the x intercept because the x value becomes 0. So any case that results in x = 0 is an intercept. When y-value is 2,3,-5 (plugged into the equation as x) we get x = 0, so we know these are the x - intercepts :) happy to clarify any of that if it wasnt clear @lolibear